# KiDS-Legacy: Redshift distributions and their calibration

### v1

@@ 2.1 Calibration data sets
The calibration sample used to estimate redshift distributions in KiDS-Legacy with the colour-based SOM method is drawn principally from the KiDZ sample described in section 5 of W24. The sample consists of 126 085126085126\,085126 085 sources drawn from $22$ spectroscopic samples/surveys, which have been compiled following a hierarchy that resolves internal and external duplicates in the data sets. The hierarchy ranks the constituents such that we keep spectra preferentially from the sample that is most likely to provide a reliable redshift. The details of this hierarchy and the homogenisation of the various redshift quality metrics are detailed in W24, the resulting redshift distribution is shown in the top panel of Fig 1.
@@
The calibration sample for clustering redshifts used in KiDS-Legacy differs from the one described in W24. As opposed to previous work (van den Busch et al., 2020; Hildebrandt et al., 2021), we only include samples that cover multiple KiDS tiles and provide sufficient contiguous overlap with KiDS or KiDZ observations, i.e. 2dFLenS (Blake et al., 2016), SDSS BOSS DR12 (LOWZ and CMASS, Alam et al., 2015), GAMA DR4 (Driver et al., 2022), and VIPERS PDR-2 (Scodeggio et al., 2018). We apply additional masking to ensure a consistent footprint between the KiDS-Legacy data, the spectroscopic data, and their provided spectroscopic random catalogues. We remove the relatively small overlap of 2dFLenS with the northern KiDS patch and limit the VIPERS data set to a redshift range of $1.180.6\leq z<$ 1.18 to be consistent with the random catalogues and to mitigate the incompleteness from the colour sampling at $z<0.6$ (Garilli et al., 2014). Finally, we add 109 381109381109\,381109 381 recently released spectra from the Dark Energy Spectroscopic Instrument (DESI Collaboration et al., 2016a, b) Early Data Release. Specifically, we use the designated clustering catalogues containing a subset of the LRG and ELG samples (see section 4.2 of DESI Collaboration et al., 2023). This new set of calibration samples for the CC method (bottom panel of Fig. 1) covers a combined total of more than 80 % of the KiDS-Legacy footprint (Fig. 2).
@@
Table 2 details all data sets utilised for calibration in KiDS-Legacy. The table indicates samples that are used for redshift calibration with the SOM calibration (Sect. 4) and those which are used for clustering redshifts (Sect. 5). Their respective redshift distributions are shown in Fig. 1.
Table 2: Spectroscopic redshift samples used for the KiDS-Legacy redshift calibration.
Survey/Field $N_{\rm spec}$ Area Density Usage
[$\text{deg}^{2}$] [$\text{arcmin}^{-2}$]
KiDZ compilation 126 085126085126\,085126 085 $19.3$ $3.77$ SOM
2dFLenS 22 6752267522\,67522 675 $382.4$ $0.02$ CC
BOSS DR12 60 4826048260\,48260 482 $422.6$ $0.04$ CC
DESI EDR 109 381109381109\,381109 381 $44.2$ $0.69$ CC
GAMA DR4 161 839161839161\,839161 839 $136.1$ $0.33$ CC
VIPERS 26 4082640826\,40826 408 $9.3$ $0.79$ CC
@@
 [^2] The KiDZ spectroscopic compilation is described in W24. VIPERS data are included in both the KiDZ compilation used by the SOM and in the sample used for cross-correlations.
@@
Figure 1: Redshift distribution of the calibration data used for KiDS-Legacy. The top panel displays the full KiDZ data in gray and the proportion of that enters each tomographic bin after calibrating the fiducial SOM as stacked histogram. The bin edges are indicated by the dashed vertical lines. The bottom panel shows the spectroscopic surveys used as calibration samples for the clustering redshift measurements (also stacked).
@@
Figure 2: Footprint of the spectroscopic surveys overlapping KiDS and used for the clustering redshift measurements. VIPERS is exclusive to the KiDZ fields, which are not shown here.
@@ 2.2 Weight assignment
One important distinction between the calibration fields and the wide-fields used for lensing is that the calibration fields lack the data-products required for lensfit shape estimation (specifically individual calibrated exposures, see W24, ). As such, sources in calibration fields that do not overlap with the wide-field data do not contain shape-measurement information, in particular the shape-measurement weights (see W24 for details about the imaging differences within the KiDZ fields). Since the shape weights correlate with the photometric observables, they present an additional selection which has to be taken into account in the SOM and the CC calibration.
Therefore, we replicate the lensfit weights in KiDZ using $k$-nearest neighbour matching. To each KiDZ galaxy we assign the lensfit weight of a galaxy from KiDS-dr5 that is closest in $r$-band magnitude (MAG_AUTO), half-light radius (FLUX_RADIUS), GAaP major-to-minor axis ratio (Bgaper / Agaper), photometric redshift (Z_B), and average PSF size per tile (PSF_RAD).
To validate this process, we run split the KiDS-Legacy wide-field sample into two halves (by splitting the survey at RA$=180$ degrees), and inherit fake lensfit weights from one half onto the other. We then compare the inherited weights to the originally measured ones. For ease of interpretation, we rescale the lensfit weights in this test to the range $w\in[0,1]$. We find that the weight inheritance is robust, having a median residual between the real and synthetic weights of precisely zero, driven by the vast majority of sources residing at a true weight of either zero or one, and being correctly assigned this limiting weight (thereby having precisely zero residual). The scatter in the weight residuals is similarly benign, at $\sigma[w_{\rm true}-w_{\rm fake}]=0.09$ (a perfectly random assignment of fake weights produces a scatter of approximately $0.6$). As such, we conclude that the weight inheritance is functioning appropriately.
@@ 2.3 Tomographic binning
A central aspect of weak lensing tomography is the choice of tomographic binning. In previous KiDS analyses, tomographic bins were defined using a set of cuts in photometric redshift ($z_{\rm B}$). For KiDS, these cuts were initially constructed to have four bins of fixed width $\Delta z_{\rm B}=0.2$ between $0.1<z_{\rm B}\leq 0.9$ (Hildebrandt et al., 2017). [^3] We emphasise here the importance of the inequalities used in these definitions: as the photo-$z$ estimates are discrete with finite steps of $0.01$, whether one uses $z_{\rm B}\leq 0.9$ or $z_{\rm B}<0.9$ has a non-negligible impact on the sample definition. With the introduction of the VIKING near-infrared data and better high-$z$ performance of the photometric redshifts, a fifth (higher redshift) tomographic bin was introduced, which used a width of $\Delta z_{\rm B}=0.3$ ($0.9<z_{\rm B}\leq 1.2$) (Hildebrandt et al., 2020). These bins resulted in tomographic bins (for the last KiDS analysis, see Hildebrandt et al. 2021) that contained between $2.8$ million (bin one) and $8.1$ million (bin three) sources.
This choice of tomography can, however, be shown to be sub-optimal for cosmic shear tomography signal-to-noise and figure-of-merit in typical applications. Sipp et al. (2021) advocate equipopulated bins as a better choice (over equidistant bins), and we opt to implement this form of tomography for KiDS-Legacy. Details of our chosen (six) tomographic bins, such as number densities and ellipticity dispersions, are provided in Table 3. It should be noted that the $z_{\rm B}$ binning is chosen a priori based on the SKiLLS simulations (see Sect. 3 and Li et al., 2022). In combination with the discreteness of $z_{\rm B}$, this leads to bins that are only approximately equipopulated in the KiDS-Legacy data.
Table 3: Properties of the six KiDS-Legacy tomographic bins and the full source sample, using our fiducial redshift calibration procedure.
Bin Selection $N$ $n_{\rm eff}$ $\sigma_{\epsilon}$ $N_{\rm gold}$ $n_{\rm eff,gold}$ $\sigma_{\epsilon,{\rm gold}}$ $m_{\rm gold}$ $n_{\rm eff,gold}/n_{\rm eff}$
@@
444 Values of $\sigma_{\epsilon}$ and $n_{\rm eff}$ are computed using Eqs. C.9 and C.12 of Joachimi et al. (2021), respectively. The $\sigma_{\epsilon}$ values correspond to the ellipticity dispersion per component. $m_{\rm gold}$ corresponds to the multiplicative shear measurement bias. Statistics in the ‘gold’ columns are computed for gold-selected sources using the gold weights described in Sect. 4.2, and include contributions from multiplicative shear biases, which are themselves given in the table.
@@ 3 Simulations
KiDS-Legacy utilises the ‘SKiLLS’ simulation of Li et al. (2022), as well as an updated version of the MICE2 simulation (Fosalba et al., 2015a, b; Crocce et al., 2015; Carretero et al., 2015) presented in van den Busch et al. (2020) and utilised in Wright et al. (2020a). SKiLLS is a multiband image simulation based on the SURFS dark-matter simulation (Elahi et al., 2018) and Shark semi-analytic model (Lagos et al., 2018), whereas MICE2 is a simulated galaxy catalogue derived from the MICE-Grand Challenge simulation, which we post-process with an analytic photometry model. Both simulations are constructed to replicate the photometric properties of KiDS and VIKING data in each of the $ugri_{1}i_{2}ZY\!J\!H\!K_{\rm s}$ bands as well as lensfit shape weights (beside other aspects like shear, clustering, etc.).
There are two additional differences between SKiLLS and MICE2 that are worthy of comment. First, MICE2 covers an on-sky area of about 5000 $\text{deg}^{2}$, whereas SKiLLS is limited to 108 $\text{deg}^{2}$. Secondly, SKiLLS has a much larger redshift baseline ($2.50.001<z<$ 2.5) than MICE2, which is limited to $0.07\lesssim z\lesssim 1.4$ and therefore does not allow us to simulate the KiDS data in the sixth tomographic bin (or possible high-$z$ tails of the other bins) with high fidelity. Due to these differences, we rely on MICE2 to simulate our clustering redshift analysis (requiring the additional area) whereas SKiLLS is our primary simulation for the SOM calibration (covering the sixth bin). Nevertheless, both simulations are useful where they overlap, since they give us additional redundancy and allow us to test how our redshift estimates depend on the assumptions underlying both simulations.
@@ 3.1 SKiLLS
The image-simulation based SKiLLS utilises imaging properties (limiting magnitudes, PSFs, etc.) sampled directly from the KiDS-$1000$ data set (Kuijken et al., 2019), such that the observational parameters are representative of the parameter distributions therein. However, the base simulations tend to overproduce sources (relative to the data) at low-resolution and high signal-to-noise, leading to possible systematic biases in the recovery of shape calibration values, and which could also cause bias in redshift distribution estimates (as the resolution and signal-to-noise ratio are correlated with colour and redshift). Therefore, in order to optimise the similarity between simulations and data, Li et al. (2022) performed an a posteriori reweighting of the simulated wide-field sources by comparing their abundance in a 2-dimensional space of shape-measurement signal-to-noise ratio and source-resolution space to the KiDS wide-field data. In KiDS-Legacy we follow the methods of Li et al. (2022), and implement a similar reweighting scheme to construct our simulated calibration samples (see Sect. 3.3) and corresponding wide-field samples (see Sect. 3.3.2).
@@ 3.2 MICE2
For KiDS-Legacy we use an updated version of the KiDS-like MICE2 mocks that resembles dr5 and implements an improved analytic photometry model (Linke et al., 2025). We derive all the necessary calibration data sets from the underlying base simulation. Previous KiDS analyses have put considerable effort into analytically mimicking their observed properties as closely as possible (van den Busch et al., 2020), by reconstructing the samples’ selection functions (typically in colour, redshift, and/or derived properties like stellar mass) in the simulation space. The documented spectroscopic success rate (as a function of redshift/magnitude) is similarly included where available. van den Busch et al. (2020) provide extensive demonstrations of the performance of this sample construction using the MICE2 simulation, which were used for the calibration of KiDS-$1000$. Generally speaking, the spec-$z$ samples were difficult to be faithfully reproduced in the simulation space without ad-hoc modifications to the original selection window, and as such were defined with a modified selection window that reproduced the expected colour, redshift, and number density distributions seen in the data.
Similar to van den Busch et al. (2020), we construct the wide-field calibration data sets such that they match observed spectroscopic data in sky coverage and relative overlap (e.g. between BOSS and GAMA), but additionally we decided to apply a stellar mask that we construct from the real KiDS-Legacy masks by tiling the MICE2 footprint. Since we use DESI and VIPERS data for the first time in a KiDS clustering redshift analysis, we implemented their respective selection functions for MICE2 similar to the existing ones for the GAMA, BOSS and 2dFLenS samples of van den Busch et al. (2020). For details refer to Appendix A.

### v2

@@ 2.1 Calibration datasets
The calibration sample used to estimate redshift distributions in KiDS-Legacy with the colour-based SOM method is drawn principally from the KiDZ sample described in Section 5 of W24. The KiDZ imaging were taken under the same observational conditions as the main KiDS and VIKING surveys and thus share the photometric properties of the wide-field dataset. The sample consists of 126 085126\,085 sources drawn from $22$ spectroscopic samples and/or surveys, which have been compiled following a hierarchy that resolves internal and external duplicates in the datasets. The hierarchy ranks the constituents such that we kept spectra preferentially from the sample that is most likely to provide a reliable redshift. The details of this hierarchy and the homogenisation of the various redshift quality metrics are detailed in W24, and the resulting redshift distribution is shown in the top panel of Fig 1.
@@
The calibration sample for clustering redshifts used in KiDS-Legacy differs from the one described in W24. As opposed to previous work (van den Busch et al., 2020; Hildebrandt et al., 2021), we only included samples that cover multiple KiDS tiles and provide sufficient contiguous overlap with KiDS or KiDZ observations, namely: the 2-degree Field Lensing Survey (2dFLenS, Blake et al., 2016), Sloan Digital Sky Survey (SDSS, DR12) Baryon Osscilation Spectroscopic Survey (BOSS) low redshift (LOWZ) and constant mass (CMASS) samples (Alam et al., 2015), Galaxy and Mass Assembly (GAMA) DR4 (Driver et al., 2022), and the VIMOS Public Extragalactic Redshift Survey (VIPERS) PDR-2 (Scodeggio et al., 2018). We applied additional masking to ensure a consistent footprint between the KiDS-Legacy data, the spectroscopic data, and their provided spectroscopic random catalogues. We removed the relatively small overlap of 2dFLenS with the northern KiDS patch and limited the VIPERS dataset to a redshift range of $0.6\leq z<1.18$ to be consistent with the random catalogues and to mitigate the incompleteness from the colour sampling at $z<0.6$ (Garilli et al., 2014). Finally, we added 109 381109\,381 recently released spectra from the Dark Energy Spectroscopic Instrument (DESI, DESI Collaboration et al., 2016a, b) Early Data Release. Specifically, we used the designated clustering catalogues containing a subset of the luminous red galaxy (LRG) and emission line galaxy (ELG) samples (see Section 4.2 of DESI Collaboration et al., 2024). This new set of calibration samples for the cross-correlation (CC) method (bottom panel of Fig. 1) covers a combined total of more than 80 % of the KiDS-Legacy footprint (Fig. 2).
@@
Table 2 details all datasets utilised for calibration in KiDS-Legacy. The table indicates samples that are used for our colour-based SOM redshift calibration (Sect. 4) and for our position-based clustering redshift calibration (Sect. 5). Their respective redshift distributions are shown in Fig. 1.
@@
Table 2: Spectroscopic redshift samples used for the KiDS-Legacy redshift calibration. [^2] The KiDZ spectroscopic compilation is described in W24. VIPERS data are included in both the KiDZ compilation used by the SOM and in the sample used for $C$s.
Survey or Field $N_{\rm spec}$ Area Density Usage
$\text{deg}^{2}$] $\text{arcmin}^{-2}$]
KiDZ compilation 126 085126\,085 $19.3$ $3.77$ SOM
2dFLenS 22 67522\,675 $382.4$ $0.02$ $C$
BOSS DR12 60 48260\,482 $422.6$ $0.04$ $C$
DESI EDR 109 381109\,381 $44.2$ $0.69$ $C$
GAMA DR4 161 839161\,839 $136.1$ $0.33$ $C$
VIPERS 26 40826\,408 $9.3$ $0.79$ $C$
@@
Figure 1: Redshift distribution of the calibration data used for KiDS-Legacy. The top panel displays the full KiDZ data in grey and the proportion of it that enters each tomographic bin after calibrating the fiducial SOM, shown as a stacked histogram. The bin edges are indicated by the vertical dashed lines. The bottom panel shows the spectroscopic surveys used as calibration samples for the clustering redshift measurements (also stacked).
@@
Figure 2: Footprint of the spectroscopic surveys overlapping KiDS used for the clustering redshift measurements. VIPERS is exclusive to the KiDZ fields, which are not shown here.
@@ 2.2 Weight assignment
One important distinction between the calibration fields and the wide-fields used for lensing is that the calibration fields lack the data-products required for lensfit shape estimation (specifically individual calibrated exposures; see W24, ). As such, sources in calibration fields that do not overlap with the wide-field data do not contain shape-measurement information, in particular the shape-measurement weights (see W24 for details about the imaging differences within the KiDZ fields). Since the shape weights correlate with the photometric observables, they present an additional selection that has to be taken into account in the SOM and the CC calibration.
Therefore, we replicated the lensfit weights in KiDZ using $k$-nearest neighbour matching. To each KiDZ galaxy we assigned the lensfit weight of a galaxy from KiDS-dr5 that is closest in $r$-band magnitude (MAG_AUTO), half-light radius (FLUX_RADIUS), GAaP major-to-minor axis ratio (Bgaper / Agaper), photometric redshift (Z_B), and average PSF size per tile (PSF_RAD). This matching process is therefore conditional on the KiDS and KiDZ data having consistent photometric properties, in particular in the $r$ band (from which all but the photometric redshift metric are defined). Fortunately, the observation of the KiDZ data was taken with the same observational requirements as the main survey data, and as such the imaging are very well matched: for example, the distribution of KiDZ $r$-band seeing and magnitude limits (which drive the statistical properties of measured sizes and fluxes) are consistent with having been drawn randomly from the parent population of KiDS $r$-band seeing and magnitude limits, as computed using the two-sample Anderson & Darling (1954) test ($p$-values of 0.66 and 0.23, respectively).
To validate the accuracy of the matching process itself, we split the KiDS-Legacy wide-field sample into two halves (by splitting the survey at RA=180=180 degrees) and inherited fake lensfit weights from one half onto the other. We then compared the inherited weights to those that were originally measured. For ease of interpretation, we rescaled the lensfit weights in this test to the range $w\in[0,1]$. We found that the weight inheritance is robust, having a median residual between the real and synthetic weights of precisely zero, being driven by the vast majority of sources residing at a true weight of either zero or one, and being correctly assigned this limiting weight (thereby having precisely zero residual). The scatter in the weight residuals is similarly benign, at $\sigma[w_{\rm true}-w_{\rm fake}]=0.09$ (a perfectly random assignment of fake weights produces a scatter of approximately $0.6$). Hence, we conclude that the weight inheritance functions appropriately.
@@ 2.3 Tomographic binning
A central aspect of weak lensing tomography is the choice of tomographic binning. In all KiDS analyses to date, tomographic bins were defined using a set of cuts in photometric redshift ($z_{\rm B}$). For the initial KiDS cosmic shear analyses, based on optical data only, these cuts were constructed to have four bins of fixed width $\Delta z_{\rm B}=0.2$ between $0.1<z_{\rm B}\leq 0.9$ (Hildebrandt et al., 2017). [^3] We emphasise here the importance of the inequalities used in these definitions: as the photo-$z$ estimates are discrete with finite steps of $0.01$, whether one uses $z_{\rm B}\leq 0.9$ or $z_{\rm B}<0.9$ has a non-negligible impact on the sample definition. With the introduction of the VIKING NIR data and better high-$z$ performance of the photometric redshifts, a fifth (higher redshift) tomographic bin was introduced, which used a width of $\Delta z_{\rm B}=0.3$ ($0.9<z_{\rm B}\leq 1.2$) (Hildebrandt et al., 2020). These selections resulted in tomographic bins that (for the last KiDS analysis, see Hildebrandt et al. 2021) contained between $2.8$ million (bin one) and $8.1$ million (bin three) sources.
This choice of tomography can, however, be shown to be sub-optimal for cosmic shear tomography signal-to-noise and figure-of-merit in typical applications. Sipp et al. (2021) advocate equipopulated bins as a better choice (over equidistant bins), and we opted to implement this form of tomography for KiDS-Legacy. Details of our chosen (six) tomographic bins, such as number densities and ellipticity dispersions, are provided in Table 3. It should be noted that the $z_{\rm B}$ binning was chosen a priori based on the SKiLLS simulations (see Sect. 3 and Li et al., 2023). In combination with the discreteness of $z_{\rm B}$, this led to bins that are only approximately equipopulated in the KiDS-Legacy data.
Table 3: Properties of the six KiDS-Legacy tomographic bins and the full source sample, using our fiducial redshift calibration procedure. [^4] Values of $\sigma_{\epsilon}$ and $n_{\rm eff}$ are computed using Eqs. C.9 and C.12 of Joachimi et al. (2021), respectively. The $\sigma_{\epsilon}$ values correspond to the ellipticity dispersion per component. $m_{\rm gold}$ corresponds to the multiplicative shear measurement bias. Statistics in the ‘gold’ columns are computed for gold-selected sources using the gold weights described in Sect. 4.2 and include contributions from multiplicative shear biases, which are themselves given in the table.
Bin Selection NN $n_{\rm eff}$ $\sigma_{\epsilon}$ $N_{\rm gold}$ $n_{\rm eff,gold}$ $\sigma_{\epsilon,{\rm gold}}$ $m_{\rm gold}$ $n_{\rm eff,gold}/n_{\rm eff}$
@@ 3 Simulations
Since the first cosmic shear analysis of KiDS (Hildebrandt et al., 2017), complex image simulations have been leveraged to calibrate shape-measurement biases (Fenech Conti et al., 2017). Subsequent analyses from Wright et al. (2020b) also utilised complex simulations to estimate informative priors on redshift distribution biases; however, these simulations were performed without an image layer. In KiDS-Legacy, we utilised the ‘SKiLLS’ suite of image simulations (Li et al., 2023) to, for the first time in KiDS, perform joint calibration of redshift distribution and shape-measurement bias parameters. Additionally, KiDS-Legacy also utilises an updated version of the MICE2 simulation (Fosalba et al., 2015a, b; Crocce et al., 2015; Carretero et al., 2015) presented in van den Busch et al. (2020) and utilised in Wright et al. (2020a). SKiLLS is a multiband image simulation based on the SURFS dark-matter simulation (Elahi et al., 2018) and Shark semi-analytic model (Lagos et al., 2018), and is constructed to match the observed multiband properties of KiDS-Legacy. SKiLLS is specifically designed to incorporate multi-wavelength imaging, realistic clustering, correlations between galaxies properties and environment, and redshift-dependent shear, thereby enabling the analysis and correction of higher-order effects in shape and redshift calibration. Conversely MICE2 is a simulated galaxy catalogue (i.e. without an image layer) derived from the MICE-Grand Challenge simulation, which we post-processed with an analytic photometry model. The two simulations were constructed to replicate the photometric properties of KiDS and VIKING data in each of the $ugri_{1}i_{2}ZY\!J\!H\!K_{\rm s}$ bands as well as lensfit shape weights (beside other aspects such as shear, clustering, etc.).
There are two additional differences between SKiLLS and MICE2 that are worthy of comment. First, MICE2 covers an on-sky area of about 5000 $\text{deg}^{2}$, whereas SKiLLS is limited to 108 $\text{deg}^{2}$. Secondly, SKiLLS has a much larger redshift baseline ($0.001<z<2.5$) than MICE2, which is limited to $0.07\lesssim z\lesssim 1.4$ and therefore did not allow us to simulate the KiDS data in the sixth tomographic bin (or possible high-$z$ tails of the other bins) with high fidelity. Additionally, as mentioned above, SKiLLS is used for joint shape-measurement calibration and redshift distribution bias estimation, which enables the correction of subtle unrecognised biases that correlate errors in shape-measurement and redshift distribution estimation (such as shear-based detection biases). Due to these differences, we relied on MICE2 to simulate our clustering redshift analysis (requiring the additional area), whereas SKiLLS was our primary simulation for the SOM calibration (covering the sixth bin). Nevertheless, the two simulations are useful where they overlap, since they give us additional redundancy and allowed us to test how our redshift estimates depend on the assumptions underlying both simulations.
@@ 3.1 SKiLLS
The multiband image-simulation based SKiLLS utilises imaging properties (limiting magnitudes, PSFs, etc.) sampled directly from the KiDS-$1000$ dataset (Kuijken et al., 2019), such that the observational parameters are representative of the parameter distributions therein. However, the base simulations tend to overproduce sources (relative to the data) at low-resolution and high signal-to-noise, leading to possible systematic biases in the recovery of shape calibration values, and which could also cause bias in redshift distribution estimates (as resolution and S/N are correlated with colour and redshift). Therefore, in order to optimise the similarity between simulations and data, Li et al. (2023) performed an a posteriori re-weighting of the simulated wide-field sources by comparing their abundance in a 2D space of shape-measurement S/N and source-resolution space to the KiDS wide-field data. In KiDS-Legacy we followed the methods of Li et al. (2023), and implemented a similar re-weighting scheme to construct our simulated calibration samples (see Sect. 3.3) and corresponding wide-field samples (see Sect. 3.3.2).
@@ 3.2 MICE2
For KiDS-Legacy we used an updated version of the KiDS-like MICE2 mocks that resembles dr5 and implements an improved analytic photometry model (Linke et al., 2025). We derived all the necessary calibration datasets from the underlying base simulation. Previous KiDS analyses have put considerable effort into analytically mimicking their observed properties as closely as possible (van den Busch et al., 2020), by reconstructing the samples’ selection functions (typically in colour, redshift, and/or derived properties such as stellar mass) in the simulation space. The documented spectroscopic success rate (as a function of redshift and magnitude) was similarly included where available. van den Busch et al. (2020) provide extensive demonstrations of the performance of this sample construction using the MICE2 simulation, which were used for the calibration of KiDS-$1000$. Generally speaking, it was difficult to faithfully reproduce the spec-$z$ samples in the simulation space without ad hoc modifications to the original selection window, and as such they were defined with a modified selection window that reproduced the expected colour, redshift, and number density distributions seen in the data.
Similar to van den Busch et al. (2020), we constructed the wide-field calibration datasets such that they match observed spectroscopic data in sky coverage and relative overlap (e.g. between BOSS and GAMA), but additionally we decided to apply a stellar mask, which we constructed from the real KiDS-Legacy masks by tiling the MICE2 footprint. Since we used DESI and VIPERS data for the first time in a KiDS clustering redshift analysis, we implemented their respective selection functions for MICE2 similar to the existing ones for the GAMA, BOSS, and 2dFLenS samples of van den Busch et al. (2020). For details refer to Appendix A.